# Data-Driven Lipschitz Continuity: A Cost-Effective Approach to Improve Adversarial Robustness

### v1

@@ 4.1 Setup
We evaluated the performance on CIFAR10, CIFAR100, and ImageNet datasets under the white-box scenario with an $L_{\infty}$ norm. To ensure comparability of results, we assessed robustness using AutoAttack [12], For CIFAR10 and CIFAR100 datasets, $\epsilon$ is set to $8/255$, while for the ImageNet dataset, $\epsilon$ is set to $4/255$. The model weights are publicly accessible from RobustBench. The ablation study involves exploring the selection of the optimal $c^{r}$, the combination of various models trained from different techniques, the verification of gradient masking, and assessments of certified adversarial robustness via randomized smoothing. Due to the page limit, the full experimental results of the ablation study are listed in Appendix.
@@ 4.2 White-box Evaluation
4.2.1 Performance Analysis on CIFAR10 and CIFAR100 Datasets
@@
(a) CIFAR10 dataset
# Method accnat accAA
* [11] + Ours 93.20 71.70
1 [48] 93.27 71.07
2 [11] 93.25 70.69
3 [49] 95.19 69.71
(b) CIFAR100 dataset
# Method accnat accAA
* [11] + Ours 74.97 44.00
1 [11] 75.22 42.67
2 [49] 83.08 41.80
3 [50] 73.85 39.18
The model used in this study is based on WRN-70-16 architecture with SiLU function while generative data were involved during the training phase. The value of $c^{\text{th}}_{i}$ is obtained by feeding all images from the training set without any augmentation, and $c^{r}$ was set to $2^{-8}$ for this experiment. Tables 1(a) and 1(b) summarize the top-3 competitors on Robustbench for CIFAR10 and CIFAR100 datasets, respectively, where ##\## represent the rankings, our results are marked by the asterisk (*), accnat and accAA denote the accuracy against clean data and adversairal examples generated by AutoAttack, respectively.
As can be seen, our method combined with WRN-70-16 with SiLU function gains improvement in robustness by at least 0.9% and achieves the best results on Robustbench for both datasets. Nevertheless, standard accuracy (accnat) is decreased. Many factors might affect the results. For example, the single additional hyper-parameter introduced in this study might not provide sufficient granularity to fit all layers in the target model.
4.2.2 Performance Analysis on ImageNet Dataset
Table 2: The results of top-3 competitors for ImageNet dataset on Robustbench.
# Method Architecutre accnat accAA
* [51] + Ours Swin-L 78.88 60.04
1 [51] Swin-L 78.92 59.56
2 [49] ConvNeXtV2-L + Swin-L 81.48 58.50
3 [51] ConvNeXt-L 78.02 58.48
In this experiment, we utilized the Swin [52] model architecture, a variant of transformers. However, scanning the approximately 1.2 million training images provided by the ImageNet dataset to determine the value of the hyper-parameter introduced in the forged function defined in (13) might take a long time. Alternatively, we randomly selected about 5,000 images as the observed images to determine the value of the hyper-parameter. Ideally, determining the optimal choice of $c^{r}$ requires conducting an ablation study to explore the relationship between the chosen $c^{r}$ and robust accuracy on a validation set. To accelerate this procedure, we first seek a value of $c^{r}$ with the highest standard accuracy. The candidate values are selected in a small range centered around this value.
Tables 2 lists the top-3 competitors on Robustbench for ImageNet dataset, including ranking, architecture, standard accuracy, and robust accuracy against AutoAttack. The experimental results demonstrate that the Swin-L model with GELU combined with our method can obtain improvements in robust accuracy and achieve the best result while standard accuracy has a tiny drop. This finding verifies that our method can be applied to both convolutional and fully connected layers.
4.2.3 Combination with Various Models
The experiment aims to assess the generability of the proposed function on adversarial trained models with identical architecture but from various training strategies and to evaluate the potential cost reduction of adversarial training. We integrated the proposed approach with partial models selected from RobustBench, whose weights are obtained directly from the official without any modifications, and also included a model trained by TRADES [53] as a baseline for CIFAR10 dataset. The selected models were trained using different techniques, such as adding perturbations in internal layers, retrieving information using knowledge distillation, reducing inefficient training data, or involving additional images from generated models or another dataset. Except for the model used in RST-WAP, which is WRN-28-10, the model architecture we utilized is WRN-34-10 with ReLU, as it is the most popular network on the RobustBench leaderboard [13].
For the white-box evaluation, the value of $c^{r}$ is set to $2^{-7}$. Table 3(a) and 3(b) present standard and robust accuracy of models integrated with our method for CIFAR10 and CIFAR100 dataset, respectively. In these tables, the column Original indicates the original results reported by RobustBench, and the column Original+Ours demonstrates the results of the proposed method. As indicated in these tables, for CIFAR10 dataset, the proposed method enhances robust accuracy by more than $2\%$ for RST-AWP, DefEAT, and LTD models, while other models receive approximately $1$ to $1.5\%$ improvement in robustness. Similarly, for CIFAR100 dataset, these models meet at least a $1\%$ increase in robustness. The empirical results prove that the resilience of existing models against adversarial attacks can be improved by Lemma 2. We believe that the proposed solution is general as it achieves great success in models incorporating different training techniques.
Another advantage of the proposed method that we would like to highlight is that the cost of our approach can almost be ignored compared to the cost of adversarial training as the cost involves only a single pass scan of a set of images to determine the hyper-parameter $c^{th}$. This implies that these models can enhance robustness for free. Specifically, LefEAT can achieve a robust accuracy of 57.30% by removing inefficient training data. By combining LefEAT model with our approach, a robust accuracy of 59.55% can be achieved, which is comparable to RST-AWP (60.04%). However, RST-AWP introduces more images from another dataset, resulting in a higher cost in each epoch. Similarly, for the CIFAR100 dataset, DefEAT with our proposed method achieves a robust accuracy of 32.11%, which is better than EffAug (31.85%), which involves more complex data augmentation during the training stage. This aligns with the suggestion by DefEAT that some data can be removed without hurting robustness. Holistically, we believe that our approach might provide a hint during the late phase of adversarial training to drop inefficient weights, resulting in further cost savings or enhanced resilience.
Another interesting observation presented in these tables is that standard accuracy is also improved for all models on both CIFAR10 and CIFAR100 datasets. Although this phenomenon cannot be explained by Lemma 2, we believe that the output range between ReLU and the proposed functions has a high similarity. Consequently, this arrangement can maintain the accuracy of clean data.
(a) CIFAR10 dataset
Method Original Original+Ours accnat accAA accnat accAA RST-AWP [54] 88.25 60.04 89.50 62.76 DefEAT [39] 86.54 57.30 87.40 59.55 LTD [55] 85.21 56.94 85.98 59.25 AWP [54] 85.36 56.17 86.19 57.85 TRADES[53] 85.34 52.86 85.78 53.80
(b) CIFAR100 dataset
Method Original Original+Ours accnat accAA accnat accAA EffAug [56] 68.75 31.85 69.14 32.57 DKLD [50] 64.08 31.65 64.26 32.58 DefEAT [39] 64.32 31.13 66.42 32.11 LTD [55] 64.07 30.59 64.29 31.95 AWP [54] 60.38 28.86 60.63 29.72
@@ 4.2.4 Gradient Masking Verification
The first item has been examined by AutoAttack as it involves three white-box attacks and one black-box attack. By compared robust accuracy shown in Table 1(a), 1(b) and 2, the models combined with the proposed method perform better robust accuracy than the original models. It indicates black-box cannot produce more adversarial examples.
The full experimental results for the rest of experiments can be found in the appendix. The results demonstrate that the proposed algorithm does not violate any of the above rules and the certified robustness improves by our method across most settings. From the evidence, we believe that the proposed method does not encounter the gradient masking problem among different hyper-parameters and various models on CIFAR10 and CIFAR100 datasets.
@@ 5 Conclusion
In this paper, we recap how robustness is certified by the theorem of Lipschitz continuity. We introduce the concept of the empirical Lipschitz constant, which minimizes the influence of the space not drawn from real data, resulting in a precise estimation of the robustness of the corresponding observed data. We prove that by remapping the input domain of a specific layer to a constrained range, the Lipschitz constant can be shrunk, leading to better robustness. The proposed function introduces only one parameter, the value of which can be determined by scanning the training data once, without re-training or fine-tuning. Compared with adversarial training, the proposed method is almost cost-free. The experimental results suggest that our method can be combined with various existing methods and achieve robustness improvements, and no gradient masking occurs in our algorithm. Furthermore, our method can achieve the best robust accuracy for CIFAR10, CIFAR100, and ImageNet datasets on the RobustBench leaderboard.
Numerous future directions merit exploration. Firstly, due to the property of maximization, the proposed function might easily be influenced by outliers. Designing a better function is an interesting research topic. Secondly, exploring the combination with various activation functions, different model architectures or large-scale datasets would be beneficial. Lastly, it is worth investigating to understand the theoretical reasons why our proposed function improves standard accuracy.
@@ A.1 Hyper-parameter Selection
The results on CIFAR10 and CIFAR100 are presented in Table 5 and Table 5, respectively. Moreover, the results of accuracy against CW attack on $L_{\infty}$ norm for CIFAR10 and CIFAR100 datasets are presented in Tables 6(a) and 6(b), respectively. As can be seen, when $c^{r}$ is set to $2^{-8}$, all models achieve better standard accuracy and robust accuracy. Additionally, the results for all models with $c^{r}=2^{-7}$ are surpassed by those when $c^{r}$ is set to $2^{-8}$. Robust accuracy can be further enhanced by setting $2^{-6}$, while standard accuracy might drop compared to the original. The results suggest that $c^{r}=2^{-7}$ is a solution that balances standard accuracy and robustness. Nevertheless, when robustness is a major concern, $c^{r}=2^{-6}$ is a better choice.
@@
Table 4: Ablation study of selecting optimal $c^{r}$ for CIFAR10 dataset.
@@
Table 5: Ablation study of selecting optimal $c^{r}$ for CIFAR100 dataset.
@@ Appendix B Full Experimental Results of Gradient Masking Verification
Table 6: The robust accuracy against CW attack on $L_{\infty}$ norm.
(a) CIFAR10 dataset
@@
Table 7(a) and 7(b) present the robust accuracy against adversarial examples generated by the original models on CIFAR10 and CIFAR100 datasets, respectively. As observed, none of the models showed lower robust accuracy than the original model. It indicates that adversarial examples can be efficiently crafted by utilizing the gradients from the victim models.
Table 8 and 9 presents the robust accuracy against FGSM and PGD attacks among different radii of the $\epsilon$-ball on the CIFAR10 and CIFAR100 datasets, respectively. As observed, the robust accuracy against FGSM, a one-step attack, is always higher than the robust accuracy against PGD, an iterative attack. This implies that the gradient is reliable, allowing the PGD attack to adjust the gradient direction multiple times to find adversarial examples. Additionally, we observe that the robust accuracy against PGD attacks for all models gradually decreases to zero as the radius of the $\epsilon$-ball increases. This indicates that the quality of gradients is preserved, enabling PGD attacks to move the gradient toward examples not in the observed distribution.

### v2

@@ 4.1 Setup
We evaluated our method on CIFAR10, CIFAR100, and ImageNet datasets under white-box attack scenarios using AutoAttack [17] with $L_{\infty}$ norm constraints ($\epsilon=8/255$ for CIFAR10/100 and $\epsilon=4/255$ for ImageNet). All baseline models were sourced from the public RobustBench repository for reproducibility.
To determine threshold values $c^{\text{th}}_{i}$, we processed the complete training sets for CIFAR10 and CIFAR100 without data augmentation. For the larger ImageNet dataset, scanning all 1.2 million training images would require substantial time. Alternatively, we randomly sampled approximately 5,000 representative images to establish thresholds efficiently. Ideally, determining the optimal choice of $c^{r}$ requires conducting an ablation study to explore the relationship between the chosen $c^{r}$ and robust accuracy on a validation set. To accelerate this procedure, we first seek a value of $c^{r}$ with the highest standard accuracy. The candidate values are selected in a small range centered around this value.
Since most models are represented in 16 bit format, and the widths of fraction bit for FP16 format defined by IEEE-754 standard and BFloat are 10 and 7 bits, respectively, truncated errors might easily occur when performing addition on two numbers with a magnitude difference of $2^{8}$ or higher. On the other hand, when $c^{r}$ is set to $2^{-5}$, all models experience a significant drop in standard accuracy, and there is meaningless in evaluating robustness at this configuration. We suggest that the candidates of $c^{r}$ are $2^{-8}$, $2^{-7}$ and $2^{-6}$ and the partial results are presented in Table 3.
Our experimental investigation included ablation studies examining: optimal scaling coefficient selection, compatibility with various adversarially trained models, verification against gradient masking, and certified robustness through randomized smoothing. Complete ablation study results are presented in Appendix A due to space constraints.
@@ 4.2 White-box Evaluation
Method Original Original+Ours
accnat accAA accnat accAA
RST-AWP [45] 88.25 60.04 89.50 62.76
DefEAT [11] 86.54 57.30 87.40 59.55
LTD [10] 85.21 56.94 85.98 59.25
AWP [45] 85.36 56.17 86.19 57.85
TRADES[51] 85.34 52.86 85.78 53.80
(a) CIFAR10 dataset
Method Original Original+Ours
accnat accAA accnat accAA
EffAug [1] 68.75 31.85 69.14 32.57
DKLD [19] 64.08 31.65 64.26 32.58
DefEAT [11] 64.32 31.13 66.42 32.11
LTD [10] 64.07 30.59 64.29 31.95
AWP [45] 60.38 28.86 60.63 29.72
(b) CIFAR100 dataset
This experiment evaluates the generalizability of the proposed method on adversarially trained models with the same architecture but trained using different strategies. Additionally, we assess the potential cost reduction from adversarial training. We integrated our approach with models selected from RobustBench, trained using various techniques such as adding perturbations in internal layers, retrieving information using knowledge distillation, reducing inefficient training data, or involving additional images from generated models or another dataset. Except for the WRN-28-10 model used in RST-WAP, all models were based on the popular WRN-34-10 with ReLU, the leading architecture on the RobustBench leaderboard [18]. The value of $c^{r}$ is set to $2^{-7}$.
Table 1(a) and 1(b) present standard and robust accuracy of models integrated with our method for CIFAR10 and CIFAR100 dataset, respectively. In these tables, accnat and accAA refer to accuracy on clean data and adversarial examples generated by AutoAttack, respectively. The column Original indicates the original results reported by RobustBench, and the column Original+Ours demonstrates the results of the proposed method. As indicated in these tables, for CIFAR10 dataset, the proposed method enhances robust accuracy by more than $2\%$ for RST-AWP, DefEAT, and LTD models, while other models receive approximately $1$ to $1.5\%$ improvement in robustness. Similarly, for CIFAR100 dataset, these models meet at least a $1\%$ increase in robustness. The empirical results prove that the resilience of existing models against adversarial attacks can be improved by Lemma 2. We believe that the proposed solution is general as it achieves great success in models incorporating different training techniques.
Additionally, models integrated with our approach incur significantly lower training costs compared to models with similar robustness. For example, combining the LefEAT model with our method achieves a robust accuracy of 59.55% can be achieved, which is comparable to RST-AWP (60.04%). However, RST-AWP introduces more images from another dataset, resulting in a higher cost in each epoch. Similarly, for the CIFAR100 dataset, DefEAT with our proposed method achieves a robust accuracy of 32.11%, which is better than EffAug (31.85%), which involves more complex data augmentation during the training stage. This result supports DefEAT’s suggestion that unnecessary data can be eliminated without sacrificing robustness. Overall, we believe that our approach offers a valuable strategy for pruning inefficient weights in the later stages of adversarial training, potentially leading to further cost savings or enhanced resilience.
An interesting observation from the tables is that standard accuracy improves for all models on both CIFAR10 and CIFAR100. While this is not directly explained by Lemma 2, we hypothesize that the output ranges of ReLU and our proposed function are similar, helping maintain accuracy on clean data.
Tables 2(a), 2(b), and 2(c) present the top-3 models on RobustBench for the CIFAR10, CIFAR100, and ImageNet datasets, respectively. The rankings are denoted by #\#, with our results marked by an asterisk (*).
For CIFAR10 and CIFAR100 datasets, we used the WRN-70-16 architecture with SiLU activations, incorporating generative data during training. As shown, our approach, when applied to the WRN-70-16 model with SiLU, improves robustness by at least 0.9%, achieving the best results on RobustBench for both datasets. However, there is a slight decrease in standard accuracy (accnat). Several factors may explain this outcome. For instance, the additional hyperparameter introduced in this study might not offer sufficient granularity to optimize all layers of the target model effectively. Although the proposed method was originally designed for ReLU activations, these results indicate its applicability to other activation functions. Future work could explore the design of more specialized functions to further optimize performance with different activation function.
On the other hand, For the ImageNet dataset, we employed the Swin model [33], a transformer-based architecture. Experimental results show that the Swin-L model with GELU activations, combined with our method, improves robust accuracy and achieves the best result, while standard accuracy sees only a marginal decrease. This suggests that our approach is effective across both convolutional and transformer-based architectures.
# Method Architecture accnat accAA
* [43] + Ours WRN-70-16 93.20 71.70
1 [37] WRN-70-16 93.27 71.07
2 [43] WRN-70-16 93.25 70.69
3 [5] WRN-70-16 95.19 69.71
(a) CIFAR10 dataset
# Method Architecture accnat accAA
* [43] + Ours WRN-70-16 74.97 44.00
1 [43] WRN-70-16 75.22 42.67
2 [5] WRN-70-16 83.08 41.80
3 [19] WRN-70-16 73.85 39.18
(b) CIFAR100 dataset
# Method Architecutre accnat accAA * [32] + Ours Swin-L 78.88 60.04 1 [32] Swin-L 78.92 59.56 2 [5] ConvNeXtV2-L + Swin-L 81.48 58.50 3 [32] ConvNeXt-L 78.02 58.48
(c) ImageNet dataset
@@
Method RobustBench $c^{r}=2^{-8}$ $c^{r}=2^{-7}$ $c^{r}=2^{-6}$
accnat accAA accnat accAA accnat accAA accnat accAA
TRADES [51] 85.34 52.86 85.57 52.97 85.78 53.80 85.49 55.37
Table 3: Ablation study of selecting optimal $c^{r}$ for CIFAR10 dataset.
4.3 Cost Analysis
For the CIFAR-10 and CIFAR-100 datasets, the total training time for AWP with WRN-34-10 on a V100 GPU is approximately 20 hours. In contrast, other methods use larger models and incorporate additional data [43, 37, 5], leading to a total training time of over 200 hours. In comparison, scanning the entire CIFAR-10 or CIFAR-100 dataset, or partial training set on ImageNet dataset, takes less than 5 minutes on a V100 GPU.
Regarding the costs of hyperparameter search, as discussed in Appendix A.1, there are only three candidate parameters to evaluate performance. We can efficiently assess white-box performance using partial data from the training set, which significantly reduces the computational time. In practice, the total time for hyperparameter search is about 1-3 hours, depending on the model size and the dataset used. This demonstrates that our approach incurs significantly lower computational overhead.
@@ 4.3.1 Gradient Masking Verification
AutoAttack has examined the first item, which involves three white-box attacks and one black-box attack. By comparing robust accuracy shown in Table 2(a), 2(b) and 2(c), the models combined with the proposed method perform better robust accuracy than the original models. It indicates black-box attacks cannot produce more adversarial examples.
The full experimental results for the rest of the experiments can be found in Appendix B. The results demonstrate that the proposed algorithm does not violate any of the above rules and the certified robustness improves by our method across most settings. From the evidence, we believe that the proposed method does not encounter the gradient masking problem among different hyper-parameters and various models on CIFAR10 and CIFAR100 datasets.
4.4 Discussion
Comparison with Lipschitz-constrained methods A common misconception is that minimizing the Lipschitz constant necessarily improves model robustness. However, in extreme cases — for example, replacing the weights of a linear layer with an identity matrix — the Lipschitz constant becomes 1, but the resulting model may fail to generalize, yielding zero natural accuracy and thus no meaningful robustness to assess. While certified training [35] and Lipschitz-constrained methods [56] can reduce the Lipschitz constant, they often introduce additional regularization terms or complex objectives, making it difficult to balance robustness and performance. In contrast, our approach enhances robustness by shrinking the space of adversarial examples, without requiring explicit minimization as in (5).
Effective adversarial training The proposed function can be constructed in a single pass over the training data, without generating adversarial examples. By suppressing values in regions associated with adversarial perturbations, our method reduces the distributional mismatch between clean and adversarial data. Although similar ideas have been explored in prior work [4], such methods typically rely on explicit adversarial training. Additionally, the empirical results show that when applied to models trained without additional generative data, our method achieves robustness comparable to or better than models that use extensive supplementary data. This implicitly suggests that unnecessary data can be eliminated without without degrading prediction quality which is aligned with previous work [11].
Structured pruning The overall procedure of the proposed algorithm shares many similarities with pruning techniques [25]. However, we would like to emphasize that the proposals are distinctly different. Pruning primarily aims to create a highly sparse model that accelerates inference times or reduces the model size for deployment on edge devices, often without considering robustness. In particular, when the pruned model exhibits extremely high sparsity without applying re-training or fine-tuning, natural accuracy drops significantly, implicitly indicating that condition (13) does not hold and that Lemma 2 is not applicable in this case. On the other hand, when the pruned model has low sparsity, only a small proportion of weights with values close to zero are eliminated, resulting in insignificant adjustments to the center and radius of the disk. The proposed algorithm, on the other hand, can identify crucial elements with minimal cost. We believe that investigating the impact of various pruning techniques, such as iterative or post-training methods, on robustness, or combining these techniques with our proposed approach, represents a valuable direction for future research.
@@ 5 Conclusion
In this paper, we revisited robustness certification through the concept of Lipschitz continuity. While existing Lipschitz-constrained methods reduce the Lipschitz constant using regularization or complex objectives, they often struggle to strike a balance between robustness and performance. In contrast, from a theoretical perspective, we demonstrated that remapping inputs to a constrained set can mitigate the influence of regions outside the real data distribution, leading to improved robustness.
Our method is data-driven and highly efficient, requiring only a single pass through the dataset to determine the appropriate parameters for constructing the forged function without fine-tuning or gradient computations. This procedure is computationally efficient and almost cost-free. The experimental results highlight that our method can be seamlessly integrated with various existing techniques to enhance robustness. When combined with models trained without additional generative data, our method achieves robustness comparable to or exceeding that of models using extensive supplementary data. These results open a promising direction for significantly reducing computational costs while maintaining or improving defensive capabilities of robust neural networks.
Numerous future directions merit exploration. For instance, investigating the integration of the proposed remapping function with different activation functions, model architectures, or large-scale datasets would be beneficial. Additionally, it would be worthwhile to explore the theoretical foundations behind the observed improvements in standard accuracy introduced by our method.
Acknowledgments
We thank the founding support from National Science and Technology Council ( 113-2622-E-007 -018 ) and the computing infrastructure provided by the National Center for High-Performance Computing, National Institutes of Applied Research (NIAR), Taiwan.
@@ A.1 Hyper-parameter Selection
The results on CIFAR10 and CIFAR100 are presented in Table 4 and Table 5, respectively. Moreover, the results of accuracy against CW attack on $L_{\infty}$ norm for CIFAR10 and CIFAR100 datasets are presented in Tables 7(a) and 7(b), respectively. As can be seen, when $c^{r}$ is set to $2^{-8}$, all models achieve better standard accuracy and robust accuracy. Additionally, the results for all models with $c^{r}=2^{-7}$ are surpassed by those when $c^{r}$ is set to $2^{-8}$. Robust accuracy can be further enhanced by setting $2^{-6}$, while standard accuracy might drop compared to the original. The results suggest that $c^{r}=2^{-7}$ is a solution that balances standard accuracy and robustness. Nevertheless, when robustness is a major concern, $c^{r}=2^{-6}$ is a better choice.
@@
Table 4: Ablation study on selecting optimal $c^{r}$ for CIFAR10 dataset.
@@
Table 5: Ablation study on selecting optimal $c^{r}$ for CIFAR100 dataset.
A.2 Scaling Training Data for Hyperparameter Tuning
This experiment investigates how the number of training images used for hyperparameter selection affects both standard accuracy and robust accuracy. As shown in Table 6, robust accuracy decreases slightly as the amount of training data increases, but it consistently remains higher than that of the competing methods.
# Method Training Data [#] accnat accAA 1 [32] + Ours 5,000 78.88 60.04 2 [32] + Ours 10,000 78.92 59.97 3 [32] + Ours 20,000 78.90 59.81 4 [32] - 78.92 59.56 5 [5] - 81.48 58.50
Table 6: Ablation study on the impact of scaling training data for ImageNet dataset.
@@ Appendix B Full Experimental Results of Gradient Masking Verification
(b) CIFAR100 dataset
Table 7: The robust accuracy against CW attack on $L_{\infty}$ norm.
Table 9(a) and 9(b) present the robust accuracy against adversarial examples generated by the original models on CIFAR10 and CIFAR100 datasets, respectively. As observed, none of the models showed lower robust accuracy than the original model. It indicates that adversarial examples can be efficiently crafted by utilizing the gradients from the victim models.
Table 10 and 11 presents the robust accuracy against FGSM and PGD attacks among different radii of the $\epsilon$-ball on the CIFAR10 and CIFAR100 datasets, respectively. As observed, the robust accuracy against FGSM, a one-step attack, is always higher than the robust accuracy against PGD, an iterative attack. This implies that the gradient is reliable, allowing the PGD attack to adjust the gradient direction multiple times to find adversarial examples. Additionally, we observe that the robust accuracy against PGD attacks for all models gradually decreases to zero as the radius of the $\epsilon$-ball increases. This indicates that the quality of gradients is preserved, enabling PGD attacks to move the gradient toward examples not in the observed distribution.
@@
Alternatively, we can empirically assess the gradient masking using Lemma 2. The first step is the same as usual: obtaining the hyperparameter $c^{\text{th}_{i}}$ for each layer. The second step involves counting the occurrences of specific elements in the input vectors that are remapped to 0. The corresponding columns of the weight matrix are then replaced with zero vectors. Table 8 presents the results, where [32] + Ours + Weight refers to the alternative implementation. As shown, the alternative implementation achieves better robust accuracy compared to the original model. However, the proposed method outperforms the alternative implementation. This is because the proposed method offers the flexibility to remap the input vectors on a sample-wise granularity, while the alternative implementation replaces the corresponding columns of the weight matrix, which has a broader influence on all samples.
Method Architecutre accnat accAA [32] + Ours Swin-L 78.88 60.04 [32] + Ours + Weight Swin-L 78.82 59.71 [32] Swin-L 78.92 59.56
Table 8: Gradient masking verification by the weight adjustment.